# Quantitative U/Th deposition and cleanliness control strategies in the JUNO site air

### v1

@@ V U/Th settling measurement
We have designed three sampling methods, namely PTFE and acrylic plate, PFA cylindrical container, and PFA flask, as shown in Figure 3. Given that containers made of PFA material typically exhibit an extremely low background, they are well-suited for ultra-low concentration U/Th measurements when combined with ICP-MS. Therefore, in this experiment, PFA flasks were used to simulate the shape of acrylic sphere, aiming to compare the U/Th deposition rates with those of other shapes. To match the sample processing containers, sampling plates were custom-made only in PTFE and acrylic materials. These were used to compare the U/Th deposition rates between acrylic and PTFE, which, to some extent, reflects the differences in deposition rates on the surfaces of PFA and acrylic glass materials.
The plate is placed in different orientations (upward, downward, and sideways) to compare the U/Th settling rates in different directions. The idea of this measurement is similar to that carried out at Pacific Northwest National Laboratory (PNNL) and the SNOLAB underground facility diVacri:2020aqc . The PFA cylindrical container simulates the settling rate on the surface of a cylindrical detector, while the PFA flask simulates the settling rate on the surface of a spherical detector, which is the form of the JUNO acrylic sphere. All the sampling plate and containers underwent the following rigorous cleaning process to minimize the external contamination:
Degreasing: all sampling plates and containers are soaked in a 0.1% Alconox solution and sonicated for 5 min, followed by rinsing with deionized water, then sonicated in deionized water for another 5 min.
Primary acid washing: the sonicated sampling plates and containers are placed in an acid tank containing 20% GINUIS nitric acid and soaked overnight.
@@
V.1 Sampling
@@
The PTFE and acrylic plates are fixed on a dust-free plastic pad with needles, with 15 plates in one group. Among them, 12 plates are stacked in pairs to simulate the sampling surfaces facing upwards and downwards, while the remaining three plates are placed sideways. This arrangement ensures that each sampling point has a sampling area of 23 cm2 in three different orientations, as shown in Figure 4.(1). Both the PFA flat-bottomed bottle and PFA flask are sampled with the bottle cap open, with both the bottle mouth and the cap facing upwards or downwards (Figure 4.(2)). To investigate whether there are differences in the settling of U/Th in open and closed spaces, we created a simple small cover and placed the PFA sampling bottles inside the cover for sampling, comparing it with the sampling bottle outside the cover (Figure 4.(2)). There are a total of four sampling points in the JUNO underground experimental halls, including the top hall, the PMT and acrylic interlayer (Figure 4.(3)), the acrylic installation platform (Figure 4.(4)), and the installation space outside the main hall. Each sampling point is equipped with three types of sampling containers and a particle size analyzer (Lighthouse Handheld 3016), which is used for continuously monitoring the particle size distribution.
V.2 Sample treatment process
@@
The acrylic sphere was assembled in the JUNO experimental hall by bonding it layer by layer. After each layer of acrylic is assembled, the seams need to be polished to improve the smoothness. During the polishing process, a large amount of acrylic powder is generated, which floats in the air and is collected by our containers. Therefore, the sample treatment process not only requires the use of acid to dissolve the inorganic salts, but also requires the removal of the acrylic powder inside through filtration or ashing methods.
@@
For the PTFE and acrylic plates, the size can fit into a 23 mL PFA cylindrical bottle, so the sampling and sample processing are both done in the PFA cylindrical bottle. The sample treatment process for all samples is as follows:
@@
Add 4 mL (20 mL) of 5% nitric acid (Fisher Scientific) and 0.5 g of hydrofluoric (HF) acid (Fisher Scientific) to a 23 mL PFA cylindrical bottle (500 mL PFA flask), and heat it for 25 minutes. The purpose of adding HF acid here is to dissolve any silicates that may be present in the sample.
@@
For the plates, after acid boiling, the plate is removed and rinsed with 1 mL of hot 5% nitric acid to ensure that any residual sample solution on the plate is also washed off.
@@
Remove organic powder: in order to remove the acrylic powder from the acid solution, the acid solution needs to be filtered or ashed. For the filtration method, all acid solutions are drawn into a syringe and filtered through a filter with a 0.2 m aperture, and heating to dry to remove HF acid (because HF acid can corrode the ICP-MS), then 2 g of 20% nitric acid is added and heated to dissolve the residual; for the ashing method, all acid solutions are transferred to a platinum crucible, placed in a muffle furnace, and completely ashed at a temperature of 600∘C, then 5 g of 20% nitric acid is added and heated to dissolve the inorganic residue in the platinum crucible. We got similar results from these two ways of treatment, so we chose filtration for future samples.
Measurement: after the acid is made up to volume, the sample is sent for ICP-MS measurement.
V.3 Blanks and recovery efficiency
@@
The blank experiment involves repeating all the pre-treatment processes for the samples, the only difference being that no sample is added. The blank results can be used to evaluate the level of external contamination in the entire pre-treatment process. The blank results of the pre-treatment process are summarized in Table 2. The PTFE is not so clean as PFA, so the blanks are above 1 pg with relatively larger standard deviation (SD). The method detection limit (MDL) at 99% confidence level (C.L.) is calculated as , where is the average blank on U/Th, -value relies on the number of replicates MDL . From the results, it can be seen that the U/Th blanks of 23 mL PFA bottle, 500 mL flask, and platinum crucible after acid boiling and filtration operations are 0.2-0.6 pg, while the blanks of PTFE and acrylic plate acid boiling and ashing are 1-2 pg.
@@
23 mL PFA bottle | acid heating+filter | 24 | 0.2 | 0.1 | 0.4 | 0.2 | 0.1 | 0.4 |
500 mL PFA flask | acid heating+filter | 16 | 0.3 | 0.1 | 0.6 | 0.5 | 0.2 | 0.9 |
Platinum crucible | acid heating+ashing | 9 | 0.4 | 0.3 | 1.3 | 0.6 | 0.3 | 1.6 |
PTFE plate (6 pieces) | acid heating+filter | 10 | 0.8 | 0.4 | 2.0 | 1.1 | 0.7 | 3.0 |
@@
The entire pre-treatment process’s U/Th recovery efficiency was determined by adding known amounts of naturally non-existent 229Th and 233U standards (preserved in 5% nitric acid) to the samples, subjecting them to the same pre-treatment process, and finally sending them for ICP-MS testing. The measurement results of 10 samples showed U/Th average recovery efficiency around 106%/108%, with a standard deviation of 5%, as shown in Figure 5. The possible reason for a recovery rate exceeding 100% is that the sampling and sample processing containers are reused in the laboratory. During the sample processing, U/Th elements can permeate into the container walls and also leach out from them. After the cleaning procedure described in Section V.2, the containers reach an equilibrium between the permeation and leaching processes during sample handling. This equilibrium is corrected for the ratio of 232Th and 238U by using the recovery rates of 229Th and 233U.
V.4 Results
Outside the JUNO underground experimental hall, there is an installation room with a cleanliness level of class 40,000-50,000. The measurement data collected from all sampling containers placed at the same location (installation room) are presented in Table 3, aiming to eliminate the interference of acrylic powder in the experimental hall. From the results in the table, the following conclusions can be drawn:
@@ V.4 Results
The settling rate of U/Th in spherical flasks (No.5) is about six times lower than in cylindrical containers (No.2);
For cylindrical bottles (No.2 and No.3) and spherical flasks (No.5 and No.6), the settling rate of U/Th with the mouth of the bottle facing upwards is 30-40 times higher than when the mouth is facing downwards.
@@
The deposition amount of U/Th on the upper surface of the PTFE plate (No.8) is similar to that on the acrylic plate (No.11), but the results show that the deposition amount on the lower surface (No.9) or side (No.10) of the PTFE plate is 4 or 5 times higher than that of the acrylic plate (No.12 and No.13). The deposition on the upper surface is a combination of dust settling and adsorption processes. From the results, the deposition on the upper surface is primarily due to gravitational settling of dust, while the deposition on the lower surface or side is mainly attributed to adsorption.
@@
1 | PFA bottle | closed | upwards | 3.0 | 3.6 |
2 | open | upwards | 28015 | 43826 | |
3 | open | downwards | 102 | 122 | |
4 | PFA flask | closed | upwards | 0.7 | 1.0 |
5 | open | upwards | 5114 | 7323 | |
6 | open | downwards | 41 | 2.00.6 | |
7 | PFA flask (1 cm) | open | upwards | 4317 | 5426 |
8 | PTFE plate | open | upwards | 1624232 | 186967 |
9 | open | downwards | 394169 | 504345 | |
10 | open | sideways | 40698 | 37125 | |
11 | Acrylic plate | open | upwards | 1166442 | 1809887 |
12 | open | downwards | 246 | 9383 | |
13 | open | sideways | 15760 | 989 |
@@
1 | Top hall | 4.8 | 69160 | 92626 | |
2 | PFA | Interlayer | 13 | 39711 | 44571 |
3 | bottle | Platform | 216 | 65772 | 80449 |
4 | Installation hall | 3.8 | 28015 | 43826 | |
5 | PFA | Top hall | 4.8 | 6743 | 7748 |
6 | flask | Interlayer | 13 | 3411 | 3419 |
7 | Installation hall | 3.8 | 5114 | 7323 |
From the results of the cylindrical container, the U/Th settling rates in the Top hall (No. 1) and the acrylic installation platform (No. 3) are similar; however, the air particle counting differ by an order of magnitude. This discrepancy is due to the fact that the acrylic installation process involves polishing the surface of the bonding area, resulting in a significant amount of acrylic powder that interferes with the measurement of airborne dust particle sizes. The U/Th settling rates in the acrylic and PMT interlayers are close to those measured outside the hall, likely because there is less human activity in this area compared to the Top hall and installation platform. For the flask container, the geometric effects lead to a settling rate that is nearly an order of magnitude lower than that of the cylindrical container.
@@ VI Application to the JUNO detector systems
We plan to evaluate the potential U/Th deposition on the surfaces of acrylic sphere, PMTs, and the SS truss during the installation process, based on the U/Th deposition rate measurements from the previous section. This will allow us to assess the impact of contamination from these detector surfaces on single-event rates in the LS after the detector is completed. As shown in Table 3, for flat plates made of PTFE and acrylic, the results for the upward-facing orientation are quite similar. Although there are differences in the results for the downward-facing and side-facing orientations, their contributions are relatively minor compared to the upward-facing results. A similar trend is observed for PFA cylindrical bottles and flasks. This indicates that gravitational settling of dust during environmental exposure plays a dominant role. Therefore, even though the previous experiments did not simulate all detector material types, we can still refer to the results from PFA cylindrical bottles and flasks when evaluating deposition rates. Since the aforementioned detector structures are uniformly spherical, we intend to use the measurement results from PFA flasks for background evaluation. Additionally, the 20 kt LS, which has been purified through a four-stage purification system and filled into acrylic sphere via a filling system, requires assessment of U/Th deposition on the surfaces of LS-contacting systems (such as pipes and containers) due to environmental exposure during installation. These systems are mostly cylindrical, so we plan to use the measurement results from PFA cylindrical containers for background evaluation. Based on the estimated U/Th deposition on detector surfaces during construction, we will further calculate the impact of these natural radioactivities on single-event rates in the LS, as well as the levels of 238U-induced radon contamination in water from water-contacting detector structures.
@@ VI.1 Acrylic inner surface
After the acrylic sphere was constructed, a 3D spray head was placed into the sphere from the top. The first step is to spray water mist to reduce the dust particle content inside the sphere, with a cleanliness improvement to class 1,000 level. The second step involves high-pressure water flushing of the inner surface of the acrylic sphere to remove the surface protective film and perform the final cleaning. After these two steps, the cleanliness levels at various heights within the sphere were measured to be in the range of class 100 to 1,000. Upon completion of the cleaning process, the upper chimney was sealed with an acrylic blind plate, and the lower chimney was fitted with a stainless steel cover to prevent extensive air exchange between the interior and exterior of the sphere. Subsequently, the remaining 8-meter section of the stainless steel chimney and the calibration room were installed. Each section was sealed with nitrogen gas after installation. Once the calibration room was fully installed, the blind plate on the upper chimney was removed, and nitrogen gas was circulated through the chimney to maintain a positive pressure within both the upper chimney and the calibration room.
@@
The total area of the acrylic inner surface amounts to 4,000 m2. The acrylic filling process involves initially filling the inside and outside of the sphere with ultra-pure water, then displacing the ultra-pure water with LS inside the sphere. The entire time for cleaning took 15 days, and after additional 15 days for chimney and pipes installation, we started the water filling of the vessel on Dec 18, 2024. Based on the experimental results from the previous section, the settling rate of U/Th in spherical flasks during exposure in a closed space and controlled environment of class 1000 is estimated to be lower than 0.02 pg/m2/d, assuming a linear change in settling rate with cleanliness and similar U/Th settling on PFA flask and acrylic sphere. Therefore, the estimated U/Th settling during the one month exposure before filling (assuming a cleanliness of class 500-5,000) is of about 1.3-13 ng. During the filling process, the acrylic sphere and water surface were also exposed to air, but the exposed acrylic surface would change with the filling process. The entire filling process took two months, and the average acrylic surface exposure is one month due to the filling process. Therefore, the estimated U/Th settling during filling period is similar to that of before filling. Assuming all U/Th settling on the surfaces of the acrylic sphere and water enters the 20 kt LS, it would contribute to 0.13-1.310-18 g/g of U/Th in LS, which is well below our requirement.
@@
One day before the completion of the acrylic sphere cleaning, approximately 6 m2 (0.15% of the whole acrylic inner surface) of protection film remained attached to the inner surface of the upper hemisphere, as checked by a person from the outer surface of acrylic sphere on Dec. 1st. This is mainly due to the epoxy adhesive at the bonding line potentially sticking to the water-soluble paper while it was not completely dry, resulting in incomplete washing away by the water. Most of the residual paper is located on the upper hemisphere of the sphere, primarily for two reasons: Firstly, during the assembly process of the lower hemisphere, the paper film at the seams is applied in reverse, meaning the adhesive side faces outward to prevent it from bonding with the epoxy. Secondly, all the seams on the lower hemisphere are removed during the final inspection process inside the sphere after its construction. It is worth noting that the amount of residue mentioned here is a rough estimate made by workers through visual inspection, and the results may be slightly exaggerated, as working on the outer surface of the sphere is relatively challenging. The radiopurity of the paper film was screened using a high-purity germanium detector, revealing activity levels of 0.5 Bq/kg for 238U (214Pb and 214Bi gammas) and 0.60.2 Bq/kg for 232Th (212Pb, 228Ac, and 208Tl gammas), based on a 300 g sample. With a density of 47 g/m2, a residual of 0.15% (equivalent to 6 m2) would result in approximately 280 g of film remaining on the inner surface of the acrylic.
@@
To assess the contamination level caused by the water-soluble paper residual upon contact with the LS, we conducted an aging experiment by soaking 25 cm2 paper film in 60 g LS at 80∘C for 24 hours calculated by the Arrhenius equation. This is equivalent to soaking at room temperature (21∘C) for 244 days. Subsequently, we measured the U/Th content in the LS before and after soaking by acid extraction and ICP-MS Li:2024gqe . The results were shown in Table 5, and we found that the U/Th leaching effect from the water-soluble paper was at a level of 200-400 pg/m2/y. Considering the 0.15% of paper residual and 20 kt LS mass, the final U/Th extraction from film to LS is 0.6-1.310-19 g/g per year, which is significantly lower than the design specifications. As for the single-event background caused by residual paper left on the acrylic inner surface, the activity is about 0.2 Bq, and this background can be effectively removed through the fiducial volume cut in the analysis.

### v2

@@ V U/Th settling measurement
We designed three types of dust-collecting containers: flat plates, a cylinder, and a sphere, as shown in Fig. 3. The flat plate was intended to examine differences in U/Th deposition rates based on orientation. The cylinder and sphere were designed to simulate geometries commonly found in our detector systems—cylinders resemble tanks often in contact with LS, while the sphere represents the structural form of the JUNO detector. By using these different shapes, we aim to better evaluate U/Th deposition rates across various types of detectors.
Given that containers made of PFA material typically exhibit an extremely low background, they are well-suited for ultra-low concentration U/Th measurements when combined with ICP-MS. To match the sample processing containers, sampling plates were custom-made only in PTFE and acrylic materials. These were used to compare the U/Th deposition rates between acrylic and PTFE, which, to some extent, reflects the differences in deposition rates on the surfaces of PFA and acrylic glass materials.
The plates were placed in different orientations (upward, downward, and sideways) to compare the U/Th settling rates in different directions. The idea of this measurement is similar to that carried out at Pacific Northwest National Laboratory and the SNOLAB underground facility di Vacri et al. (2021). The PFA bottle simulates the settling rate on the surface of a cylindrical detector, while the PFA flask simulates the settling rate on the surface of a spherical detector, which is the form of the JUNO acrylic sphere.
V.1 Plates and containers cleaning
All the sampling plate and containers underwent the following rigorous cleaning process to minimize the external contamination:
Degreasing: all sampling plates and containers are soaked in a 0.1% Alconox (Alconox, Inc.) solution and sonicated for 5 min, followed by rinsing with deionized water (produced by the Milli-Q equipment), then sonicated in deionized water for another 5 min.
Primary acid washing: the sonicated sampling plates and containers are placed in an acid tank containing 20% nitric acid (GINUIS) and soaked overnight.
@@ V.1 Plates and containers cleaning
The sampling plates and containers will then undergo a blank test, as detailed in Section V.4. They are approved for use only if the 238U/232Th concentrations measured in the acid solution are consistent with the average blank value.
V.2 Sampling
Each plate has a single-side surface area of approximately 4 cm2, so six such sides give a total surface area of 23 cm2. To sample these surfaces in different orientations with the same sampling area, we use the following approach:
For side sampling, a thin needle is used to pierce the sheet and secure it onto a dust-free plastic pad, allowing both sides to collect dust. Thus, only three sheets are needed for one sample in this orientation.
For upward- and downward-facing sampling, to prevent the non-sampling side from collecting dust, two sheets are stacked together and fixed with a thin needle onto a dust-free plastic pad. This means six sheets are required for upward-facing and six for downward-facing sampling, totaling twelve sheets.
Therefore, a full set of samples requires 15 sheets in total. For each sampling location, we collect 2–3 parallel samples. Figure 4.(1) shows two parallel samples for each of the two types of sampling sheets.
Both the PFA bottle and PFA flask are sampled with the caps kept open, with both the bottle mouth and the cap facing upwards or downwards (Fig. 4.(2)). To investigate whether there are differences in the settling of U/Th in open and closed spaces, we created a simple small cover and placed the PFA sampling bottles inside the cover for sampling, comparing it with the sampling bottle outside the cover (Fig. 4.(2)).
@@ V.2 Sampling
There are a total of four sampling points in the JUNO underground experimental halls, including the Top Hall, the PMTs and acrylic interlayer (Fig. 4.(3)), the acrylic installation platform (Fig. 4.(4)), and the installation space outside the main hall. Each sampling point is equipped with three types of sampling containers and a particle size analyzer (Lighthouse Handheld 3016), which is used for continuously monitoring the particle size distribution.
V.3 Sample treatment process
@@
The acrylic sphere was assembled in the JUNO experimental hall by bonding it layer by layer. After each layer of acrylic is assembled, the seams need to be polished to improve the smoothness. During the polishing process, a large amount of acrylic powder is generated, which floats in the air and is collected in our containers. Therefore, the sample treatment process not only requires the use of acid to dissolve the inorganic salts, but also requires the removal of the acrylic powder inside through filtration or ashing methods.
@@
Both the PTFE and acrylic plates can fit into a 23 mL PFA bottle, so the sampling and sample processing are both done in the PFA bottle. The sample treatment process for all samples is as follows:
@@
Add 4 mL (20 mL) of 5% nitric acid (Fisher Scientific) and 0.5 g of hydrofluoric (HF) acid (Fisher Scientific) to a PFA bottle (PFA flask), and boil it for 25 min. The purpose of adding HF acid here is to dissolve any silicates that may be present in the sample.
@@
For the plates, after acid boiling, the plate is removed and rinsed with 1 mL of boiled 5% nitric acid to ensure that any residual sample solution on the plate is also washed off.
@@
Remove organic powder: in order to remove the acrylic powder from the acid solution, the acid solution needs to be either filtered or ashed. For the filtration method, all acid solutions are collected using a syringe and filtered through a membrane with a pore size of 0.2 m. The filtered liquid is subsequently evaporated to dryness to eliminate HF acid, which could otherwise corrode the ICP-MS. Thereafter, 2 g of 20% nitric acid is added, and the mixture is boiled to dissolve the residue. For the ashing method, all acid solutions are transferred to a platinum crucible and placed in a muffle furnace, where they are completely ashed at 600∘C. Then, 5 g of 20% nitric acid is added and boiled to dissolve the inorganic residue retained in the crucible. We got similar results from these two methods of treatment, so we chose filtration for future samples.
Measurement: after the acid is made up to volume (2 mL), the sample is sent for ICP-MS measurement.
V.4 Blanks and recovery efficiency
@@
The blank experiment involves repeating all the pre-treatment processes for the samples, the only difference being that no sample is added. The blank results can be used to assess the level of external contamination introduced during the entire pre-treatment process, including contributions from the vessel surfaces and ambient air. Subsequently, a fixed volume of 2 mL of the acid solution is transferred to the ICP-MS for analysis. By multiplying the measured concentration obtained from ICP-MS by the 2 mL solution volume, the exact mass of contamination attributable to that vessel and procedural step can be determined. The blank results of the pre-treatment process are summarized in Table 1. The PTFE is not as clean as PFA, so the blanks are above 1 pg with relatively larger standard deviation (SD). The method detection limit (MDL) at 99% confidence level (C.L.) is calculated as , where is the average blank on U/Th, -value relies on the number of replicates 821-R-16-006 . From the results, it can be seen that the U/Th blanks of PFA bottle, PFA flask, and platinum crucible after acid boiling and filtration operations are 0.2-0.6 pg, while the blanks of PTFE plate and acrylic plate acid boiling and ashing are 1-2 pg.
@@
| PFA bottle | acid boiling+filter | 24 | 0.2 | 0.1 | 0.4 | 0.2 | 0.1 | 0.4 |
| PFA flask | acid boiling+filter | 16 | 0.3 | 0.1 | 0.6 | 0.5 | 0.2 | 0.9 |
| Platinum crucible | acid boiling+ashing | 9 | 0.4 | 0.3 | 1.3 | 0.6 | 0.3 | 1.6 |
| PTFE plate (6 pieces) | acid boiling+filter | 10 | 0.8 | 0.4 | 2.0 | 1.1 | 0.7 | 3.0 |
@@
All sampling plates and containers must undergo this blank test. Only those that yield normal results are approved for use in sampling.
The entire pre-treatment process’s U/Th recovery efficiency was determined by adding known amounts of naturally non-existent 229Th and 233U standards (preserved in 5% nitric acid) to the samples, subjecting them to the same pre-treatment process, and finally sending them for ICP-MS testing. The measurement results of 10 samples showed U/Th average recovery efficiency around 106%/108%, with a standard deviation of 5%, as shown in Fig. 5. The possible reason for a recovery rate exceeding 100% is that the sampling and sample processing containers are reused in the laboratory. During the sample processing, U/Th elements can permeate into the container walls and also leach out from them. After the cleaning procedure described in Section V.3, the containers reach an equilibrium between the permeation and leaching processes during sample handling. This equilibrium is corrected for the ratio of 232Th and 238U by using the recovery rates of 229Th and 233U.
V.5 Results
Outside the JUNO underground experimental hall, there is an installation room with a cleanliness level of class 40,000-50,000. The measurement data collected from all sampling containers placed at the same location (installation room) are presented in Table 2, aiming to eliminate the interference of acrylic powder in the experimental hall. For samples, the U/Th concentration in the 2 mL solution is similarly measured via ICP-MS. Multiplying this concentration by the 2 mL volume yields the total mass of U/Th present in the solution. Subtracting the corresponding blank mass (from Table 1) gives the net deposited mass of U/Th originating from the sample itself. This net mass is then divided by the sample’s surface area and exposure duration to calculate the deposition rate (pg/d/m2).
From the results in the table, the following conclusions can be drawn:
@@ V.5 Results
The settling rate of U/Th in spherical flasks (No.5) is about six times lower than in cylindrical bottles (No.2);
For cylindrical bottles (No.2 and No.3) and spherical flasks (No.5 and No.6), the settling rate of U/Th with the mouth of the bottle facing upwards is 10-40 times higher than when the mouth is facing downwards.
@@
The deposition amount of U/Th on the upper surface of the PTFE plate (No.8) is similar to that on the acrylic plate (No.11), but the results show that the deposition amount on the lower surface (No.9) or side (No.10) of the PTFE plate is 3-20 times higher than that of the acrylic plate (No.12 and No.13). The deposition on the upper surface is a combination of dust settling and adsorption processes. From the results, the deposition on the upper surface is primarily due to gravitational settling of dust, while the deposition on the lower surface or side is mainly attributed to adsorption.
@@
| 1 | PFA bottle | closed | upwards | 3.0 | 3.6 |
| 2 | open | upwards | 28015 | 43826 | |
| 3 | open | downwards | 102 | 122 | |
| 4 | PFA flask | closed | upwards | 0.7 | 1.0 |
| 5 | open | upwards | 5114 | 7323 | |
| 6 | open | downwards | 41 | 2.00.6 | |
| 7 | PFA flask (1 cm) | open | upwards | 4317 | 5426 |
| 8 | PTFE plate | open | upwards | 1624232 | 186967 |
| 9 | open | downwards | 394169 | 504345 | |
| 10 | open | sideways | 40698 | 37125 | |
| 11 | Acrylic plate | open | upwards | 1166442 | 1809887 |
| 12 | open | downwards | 246 | 9383 | |
| 13 | open | sideways | 15760 | 989 | |
@@
| 1 | Top Hall | 4.8 | 69160 | 92626 | |
| 2 | PFA | Interlayer | 13 | 39711 | 44571 |
| 3 | bottle | Platform | 216 | 65772 | 80449 |
| 4 | Installation hall | 3.8 | 28015 | 43826 | |
| 5 | PFA | Top Hall | 4.8 | 6743 | 7748 |
| 6 | flask | Interlayer | 13 | 3411 | 3419 |
| 7 | Installation hall | 3.8 | 5114 | 7323 | |
From the results of the PFA bottle, the U/Th settling rates in the Top Hall (No. 1) and the acrylic installation platform (No. 3) are similar; however, the air particle counting differs by an order of magnitude. This discrepancy is due to the fact that the acrylic installation process involves polishing the surface of the bonding area, resulting in a significant amount of acrylic powder that interferes with the measurement of airborne dust particle sizes. The U/Th settling rates in the acrylic and PMTs interlayers are close to those measured outside the hall, likely because there is less human activity in this area compared to the Top Hall and installation platform. For the flask container, the geometric effects lead to a settling rate that is nearly an order of magnitude lower than that of the PFA bottle.
@@ VI Application to the JUNO detector systems
Background contributions to the JUNO detector from operational and environmental contamination during installation may arise through the following mechanisms:
The inner surface of the acrylic vessel was protected by film throughout most of the installation process. However, in some areas, the film may have detached, exposing the surface to air and resulting in deposition of natural radionuclides. Furthermore, during the final flushing process, residual adhesive film and particulate impurities that were not completely removed could contaminate the LS.
The purification and filling systems associated with the LS involve extensive use of pipelines, tanks, connectors, and instrumentation. These components underwent dedicated cleaning and inspection procedures to ensure surface cleanliness. Similar to the acrylic cleaning process, residual impurities that were not fully removed could potentially contaminate the LS. Furthermore, after cleaning, exposure to air during installation resulted in the deposition of additional impurities on the surfaces. During LS production and transfer, these contaminants may be flushed into the LS.
Detector components located outside the acrylic vessel that do not directly contact the LS—such as structural supports and external modules—may still contribute to the background. Although they do not directly contaminate the LS, gamma rays emitted from natural radionuclides within these materials can penetrate the water shielding and the acrylic vessel, depositing energy in the LS.
The following sections will quantitatively assess the impact of each of these three pathways based on available measurement data.
@@ VI.1 Acrylic inner surface
After the acrylic sphere was constructed, a 3D spray head was placed into the sphere from the top. The first step is to spray UPW mist to reduce the dust particle content inside the sphere, with a cleanliness improvement to class 1,000 level. The second step involves high-pressure UPW flushing of the inner surface of the acrylic sphere to remove the surface protective film and perform the final cleaning. It is challenging to directly measure the level of contaminants remaining on the acrylic surface after flushing. Instead, the effectiveness of the cleaning process was evaluated by monitoring the particulate content in water samples discharged from the bottom outlet. Flushing was continued until the particle concentration in the water consistently met the Level 50 standard (MIL-STD-1246C)111Cleanliness Level 50 standard of MIL-STD-1246C specifies a maximum particle count per liter of: 530 for particles 5 m, 230 for particles 15 m, 34 for particles 25 m, and 10 for particles 50 m. and stabilized. As a result, the amount of residual contaminants that were not removed from the acrylic surface remains undetermined. The actual cleanliness of the acrylic surface can only be verified through experimental data. After completing these two steps, the cleanliness level inside the acrylic sphere was measured at various heights and found to range between Class 100 and Class 1,000.
Upon completion of the cleaning process, the upper chimney was sealed with an acrylic blind plate, and the lower chimney was fitted with a SS cover to prevent extensive air exchange between the interior and exterior of the sphere. Subsequently, the remaining 8-meter section of the SS chimney and the calibration room were installed. Each section was sealed with nitrogen gas after installation. Once the calibration room was fully installed, the blind plate on the upper chimney was removed, and nitrogen gas was circulated through the chimney to maintain a positive pressure within both the upper chimney and the calibration room.
@@
The total area of the acrylic inner surface amounts to 4,000 m2. The acrylic filling process involves initially filling the inside and outside of the sphere with UPW, then displacing the UPW with LS inside the sphere. The entire time for cleaning took 15 days, and after additional 15 days for chimney and pipes installation, we started the UPW filling of the vessel on Dec 18, 2024. Based on the experimental results from the previous section, the settling rate of U/Th in spherical flasks during exposure in a closed space and controlled environment of class 1,000 is estimated to be lower than 0.02 pg/m2/d, assuming a linear change in settling rate with cleanliness and similar U/Th settling on PFA flask and acrylic sphere. Therefore, the estimated mass of U/Th settling on the entire acrylic inner surface during the one-month pre-filling exposure period (assuming a cleanliness level of Class 500-5,000) is approximately 1.3 to 13 ng. During the filling process, the acrylic sphere and water surface were also exposed to air, but the exposed acrylic surface would change with the filling process. The entire filling process took two months, and the average acrylic surface exposure is one month due to the filling process. Therefore, the estimated U/Th settling during filling period is similar to that of before filling. Assuming all U/Th settling on the surfaces of the acrylic sphere and water enters the 20 kt LS, it would contribute to 0.13-1.310-18 g/g of U/Th in LS, which is well below our requirement.
@@
One day before the completion of the acrylic sphere cleaning, approximately 6 m2 (0.15% of the whole acrylic inner surface) of protection film remained attached to the inner surface of the upper hemisphere, as checked by a person from the outer surface of acrylic sphere on Dec. 1st. This is mainly due to the epoxy adhesive at the bonding line potentially sticking to the water-soluble paper while it was not completely dry, resulting in incomplete washing away by the water. Most of the residual paper is located on the upper hemisphere of the sphere, primarily for two reasons: Firstly, during the assembly process of the lower hemisphere, the paper film at the seams is applied in reverse, meaning the adhesive side faces outward to prevent it from bonding with the epoxy. Secondly, all the seams on the lower hemisphere are removed during the final inspection process inside the sphere after its construction. It is worth noting that the amount of residue mentioned here is a rough estimate made by workers through visual inspection, and the results may be slightly exaggerated, as working on the outer surface of the sphere is relatively challenging. The radiopurity of the paper film was screened using a high-purity germanium detector, revealing activity levels of 0.5 Bq/kg for 238U (214Pb and 214Bi gammas) and 0.60.2 Bq/kg for 232Th (212Pb, 228Ac, and 208Tl gammas) by assuming the chain is at an equilibrium state, based on a 300 g sample. With a density of 47 g/m2, a residual of 0.15% (equivalent to 6 m2) would result in approximately 280 g of film remaining on the inner surface of the acrylic.
@@
To assess the contamination level caused by the water-soluble paper residual upon contact with the LS, we conducted an aging experiment by soaking 25 cm2 paper film in 60 g LS at 80∘C for 24 hours. For the aging test, a 10∘C increase in temperature yields an acceleration factor of approximately 2.5 International Electrotechnical Commission (2017), with the activation energy parameter set to about 80 kJ/mol. Consequently, soaking the paper film in LS at 80∘C for 24 hours is equivalent to aging at room temperature (21∘C) for approximately 244 days. Subsequently, we measured the U/Th content in the LS before and after soaking by acid extraction and ICP-MS Li et al. (2025). The results were shown in Table 4, and we found that the U/Th leaching effect from the water-soluble paper was at a level of 200-400 pg/m2/y. Considering the 0.15% of paper residual and 20 kt LS mass, the final U/Th extraction from film to LS is 0.6-1.310-19 g/g per year, which is significantly lower than the design specifications and expected runtime of JUNO. As for the single-event background caused by residual paper left on the acrylic inner surface, the measured activity of the 280 g residual film is approximately 0.2 Bq, with a U/Th concentration level of 0.5–0.6 Bq/kg. This background can be effectively removed through the fiducial volume cut in the analysis.